Calculate the 20th, 50th, and 80th percentiles for the following data set: (Use the methodology shown in Section 3.2 of the text to calculate the percentiles. Do not round intermediate calculations. Round your final answers to 1 decimal place.)

125 221 203 348 240 212 317
Answer #1

Solution:
Arranging Observations in the ascending order, We get :
125,203,212,221,240,317,348
Here, n = 7

P20 = value of (20 (n + 1 ) / 100 )th value of Observations

   =value of ((20 * 8 ) / 100 )th value of Observations

= value of 1.6th value of Observations

= 1st Observations + 0.6 ( 2nd - 1st )

= 125 + 0.6 (203 - 125 )

= 125 + 0.6 ( 78 )

= 125 + 46.8

= 171.8

P20 == 171.8

P50 = value of (50 (n + 1 ) / 100 )th value of Observations

   =value of ((50 * 8 ) / 100 )th value of Observations

= value of 4th value of Observations

= 221

P50 = 221

P80 = value of (80 (n + 1 ) / 100 )th value of Observations

   =value of ((80 * 8 ) / 100 )th value of Observations

= value of 6.04th value of Observations

= 6th Observations + 0.4 ( 7th - 6th )

= 317 + 0.4 (348- 317 )

= 317 + 0.4 ( 31 )

= 317 + 12.4

= 329.4

P80 = 329.4